A warehouse worker is pushing a 90.0 kg crate with a horizontal force of 292 N at a speed of v = 0.875 m/s across the warehouse floor. He encounters a rough horizontal section of the floor that is 0.75 m long and where the coefficient of kinetic friction between the crate and floor is 0.359.

(a)

Determine the magnitude and direction of the net force acting on the crate while it is pushed over the rough section of the floor.

magnitude N direction

(b)

Determine the net work done on the crate while it is pushed over the rough section of the floor.

J

(c)

Find the speed of the crate when it reaches the end of the rough surface.

m/sm
Answer #1

Magic Dato Aus. d fix= = M N = Ming fis (a) 292 292 - 01359 x 90 x 918 umg = -24.638 Iner = Ay. (6) a = Fret - co zitu t 8.0
Answer #2

SOLUTION :


a.


N = mg = 90*9.8 = 882 (N)


F = µk * N = 0.359*882 = 316.64 (N)


Net Push force on the crate = P - F = 292 - 316.64 = - 24.64 (N) 


So, 


Magnitude = 24.64 (N) and direction : opposite to motion.

(ANSWER)


b.


WD = | P net * d | = | - 24.64*0.75 | = 18.48 J (ANSWER).


c.


Let v’ m/s be the speed at the end of rough surface


So,


Change of K. E. = WD

=> 1/2 m (v’^2  - v^2) = - 18.48

=> 1/2 * 90 (v’^2 - 0.875^2) = - 18.48

=> v’^2 = - 18.48/45 + 0.875^2 = 0.355

=> v’ = sqrt(0.355) = 0.596 m/s (ANSWER)

answered by: Tulsiram Garg